### v1

# Averaging of quantum channels via channel-state duality

Abstract
Twirling, uniform averaging over symmetry actions, is a standard tool for reducing the description of quantum states and channels to symmetry-invariant data. We develop a framework for averaging quantum channels based on channel-state duality that converts pre- and post-processing averages into a group twirl acting directly on the Choi operator. For arbitrary unitary representations on the input and output spaces, the twirled channel is obtained as an explicit projection onto the commutant of the induced representation on $\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in}$. In the collective setting, where the commutant is the walled Brauer algebra, we introduce a partial-transpose reduction that maps channel twirling to an ordinary Schur-Weyl twirl of the partially transposed Choi operator, enabling formulas in terms of permutation operators. We further extend the construction beyond compact symmetries to reductive non-unitary groups via Cartan decomposition, yielding a weighted sum of invariant-sector projections with weights determined by the Abelian component. Finally, we provide two finite realizations of channel averaging. The first one is a “dual” averaging protocol as a convex mixture of unitary-$1$-design channels on invariant sectors. The second one is a notion of channel $t$-designs induced by weighted group $t$-designs for $t=t_{\rm in}+t_{\rm out}$.
1 Introduction
Following ideas of [1, 2, 3, 4] we define twirling of quantum channels as follows. Firstly let us introduce notation inspired by previous works on averaging of quantum states [5, 6]. In this work we assume that the dimensions of elementary systems constituting input and output subspaces are equal, however the number of elementary subsystems can be different for input and output spaces. Hence the input and output Hilbert spaces are specifed respectively as: $\mathcal{H}_{\operatorname{in}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{in}}}$ and $\mathcal{H}_{\operatorname{out}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{out}}}$. We consider arbitrary channels $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ mapping $t_{\operatorname{in}}$ qu-$d$-its to $t_{\operatorname{out}}$ qu-$d$-its.
Following [6] let us introduce two arbitrary unitary representations $\pi_{\operatorname{U}}^{\operatorname{in}}$ and $\pi_{\operatorname{U}}^{\operatorname{out}}$ of a unitary group $\operatorname{U}(d)$, acting respectively on $\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ and $\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ via maps:
$\mathcal{T}_{U}^{\operatorname{in}}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{\operatorname{U}}(U)\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger},\,\,\sigma_{\operatorname{in}}\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$
$\mathcal{T}_{U}^{\operatorname{out}}(\sigma_{\operatorname{out}})$ $=$ $\pi^{\operatorname{out}}_{\operatorname{U}}(U)\sigma_{\operatorname{out}}\pi^{\operatorname{out}}_{\operatorname{U}}(U)^{\dagger},\,\,\sigma_{\operatorname{out}}\in\mathcal{B}(\mathcal{H}_{\operatorname{out}}).$ (1)
Clearly we have: $\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\sigma_{\operatorname{in}})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)$, since $\pi^{\operatorname{in}}_{\operatorname{U}}(U^{-1})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{-1}=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}$. The last property holds because any representation is a group homomorphism. Finally, let us define twirling of quantum channels as follows (see [1], Sec. II C):
$\mathcal{M}^{\Phi}_{U}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\left(\Phi\left(\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)\right)\right),$ (2)
in which the integration is performed with respect to a Haar measure on the unitary group.
Starting from the above definition of channel twirling as uniform pre- and post-processing by a symmetry representation, we derive an explicit Choi-level expression showing that twirling of channels is equivalent to a group twirl acting directly on the Choi operator under the induced representation $\pi^{\rm out}_{\rm U}\otimes\bar{\pi}^{\rm in}_{\rm U}$, and we express the result as a projection onto the commutant algebra via generalized Schur–Weyl duality. In the collective setting with $t_{\rm in}$ input and $t_{\rm out}$ output subsystems, where the relevant commutant is the walled Brauer algebra, we introduce a partial-transpose reduction that removes the contragredient action and maps the channel twirl to an ordinary Schur-Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\rm in}+t_{\rm out})}$, enabling practical formulas in terms of permutation operators and avoiding an explicit construction of walled Brauer idempotents. Beyond compact symmetries, we extend the approach to averaging over non-unitary operations described by reductive groups using Cartan decomposition and normalized non-compact components, which yields a weighted decomposition of the twirl into invariant-sector projections with coefficients determined solely by the Abelian Cartan part. We also provide two finite realizations that parallel the role of designs in sampling: first, a dual averaging representation in which the (Cartan) channel twirl is implemented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors via embedded unitary operator bases, and second, a design-like reconstruction theorem for collective actions showing that any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$ induces a channel $t$-design that reproduces the twirled Choi operator from a finite sum.
Main results
Theorem 1 (Channel twirl in Choi form, Section 2).
Let $\Phi:\mathcal{B}(\mathcal{H}_{\rm in})\to\mathcal{B}(\mathcal{H}_{\rm out})$ be a quantum channel with Choi operator $J_{\Phi}=(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\rm{in}})(\rho_{\rm ME})\in\mathcal{B}(\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in})$. Given unitary representations $\pi^{\rm in}_{\rm U}$ on $\mathcal{H}_{\rm in}$ and $\pi^{\rm out}_{\rm U}$ on $\mathcal{H}_{\rm out}$ representing the symmetry transformations on input and output subsystems, define the twirl
$M_{U}^{\Phi}(\rho)=\int_{\mathrm{U}(d)}dU\;T^{\rm out}_{U}\circ\Phi\circ(T^{\rm in}_{U^{-1}})(\rho),\quad T^{\rm in/out}_{U}(X)=\pi_{\rm U}^{\rm in/out}(U)\,X\,\pi_{\rm U}^{\rm in/out}(U)^{\dagger}.$
Then the Choi operator of the twirled channel satisfies
$J_{M_{U}^{\Phi}}=\int_{\mathrm{U}(d)}dU\;\Big(\pi_{\rm U}^{\rm out}(U)\otimes\bar{\pi}_{\rm U}^{\rm in}(U)\Big)\,J_{\Phi}\,\Big(\pi_{\rm U}^{\rm out}(U)\otimes\bar{\pi}_{\rm U}^{\rm in}(U)\Big)^{\dagger},$
i.e. channel twirling is a state twirl on the Choi operator under $\pi_{\rm U}^{\rm out}\otimes\bar{\pi}_{\rm U}^{\rm in}$.
Corollary 1 (Collective unitary twirl reduces to ordinary Schur–Weyl twirl, Section 2).
Assume collective representations $\pi_{\rm U}^{\rm in}(U)=U^{\otimes t_{\rm in}}$ and $\pi_{\rm U}^{\rm out}(U)=U^{\otimes t_{\rm out}}$. Let $\Gamma$ denote partial transpose on the input factor of $\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in}$. Then
$J_{M_{U}^{\Phi}}=\left[\int_{\mathrm{U}(d)}dU\;\Big(U^{\otimes(t_{\rm in}+t_{\rm out})}\Big)\,J_{\Phi}^{\Gamma}\,\Big(U^{\otimes(t_{\rm in}+t_{\rm out})}\Big)^{\dagger}\right]^{\Gamma}.$
In particular, evaluating $M_{U}^{\Phi}$ can be reduced to a standard Schur–Weyl projection of $J_{\Phi}^{\Gamma}$, so one may work directly with permutation operators on $(t_{\rm in}+t_{\rm out})$ tensor factors.
Theorem 2 (Reductive (Cartan) twirl and invariant-sector decomposition, Section 3).
Let $\mathcal{G}$ be a reductive group with Cartan decomposition $G=KAK^{\prime}$, where $G\in\mathcal{G}$, $K\in\mathcal{K}$ and $A\in\mathcal{A}$, $\mathcal{K}$ denotes the maximal compact subgroup of $\mathcal{G}$, $\mathcal{A}$ denotes the maximal Abelian subgroup, and let $\pi_{\mathcal{G}}^{\rm in/out}$ be finite-dimensional matrix representations of $\mathcal{G}$ on $\mathcal{H}_{\rm in/out}$. Define the (normalized) twirl by integrating over $\mathcal{K}\times\mathcal{A}\times\mathcal{K}$ and using the adjoint action on the input side. Then the Choi operator of the twirled channel admits an invariant-sector decomposition
$J_{M_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\mathcal{P}_{k}(J_{\Phi}),$
where $\{\mathcal{P}_{k}\}$ are the commutant-sector projections determined by the restriction of $\pi_{\mathcal{G}}^{\rm out}\otimes\bar{\pi}_{\mathcal{G}}^{\rm in}$ to the maximal compact subgroup $\mathcal{K}$, while the coefficients $\beta_{k}$ depend only on the Abelian component $\mathcal{A}$ through a real-valued integral. Moreover, in the collective case (e.g. $\mathcal{G}=\mathrm{SL}(d,\mathbb{C})$ with $\mathcal{K}=\mathrm{SU}(d)$ and ${\pi}_{\mathcal{G}}^{\rm in/out}=G^{\otimes t_{\rm in}/t_{\rm out}}$) one may again express $\mathcal{P}_{k}$ using ordinary Schur–Weyl permutation operators after partial transpose.
Theorem 3 (Dual averaging as a mixture of unitary-$1$-design channels, Section 4).
The Cartan twirl $M_{\mathcal{G}}^{\Phi}$ of the Choi operator can be represented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors:
$J_{M_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\frac{1}{(D_{k}^{G})^{2}}\sum_{l=1}^{(D_{k}^{G})^{2}}\tilde{\gamma}^{(k)}_{l}\,J_{\Phi}\,\tilde{\gamma}^{(k)\dagger}_{l}$,
where $\{\tilde{\gamma}^{(k)}_{l}\}$ are Kraus operators obtained by embedding a unitary operator basis on $\mathbb{C}^{D_{k}^{G}}$ into the full representation space via the duality basis operators.
Proposition 1 (Channel $t$-design induced by a weighted group $t$-design, Section 5).
Let $t=t_{\rm in}+t_{\rm out}$ and suppose $\{(G_{i},w_{i})\}$ is a weighted group $t$-design for a subgroup $\mathcal{G}\subseteq\mathrm{GL}(d,\mathbb{C})$. Then the averaged Choi operator can be reconstructed from a finite sum:
$J_{M_{\mathcal{G}}^{\Phi}}=\left[\int_{\mathcal{G}}dG\;G^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G^{\otimes t})^{\dagger}\right]^{\Gamma}=\left[\sum_{i}w_{i}\;G_{i}^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G_{i}^{\otimes t})^{\dagger}\right]^{\Gamma}.$
2 Averaging quantum channels over arbitrary unitary transformations
In this section we provide an analytic formula for the action of the channel (2) utilizing channel-state duality. Let $J_{\Phi}=(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left(\rho_{\operatorname{ME}}\right)\in\mathcal{B}(\mathcal{H}_{\operatorname{out}})\otimes\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ denote the Choi matrix of the channel $\Phi$, in which $\rho_{\operatorname{ME}}$ is a non-normalized density matrix of a maximally entangled state on $\mathcal{H}_{\operatorname{in}}\otimes\mathcal{H}_{\operatorname{in}}$ and ${\rm 1\kern-2.20001ptl}_{\operatorname{in}}$ is the identity channel on $\mathcal{B}(\mathcal{H}_{\operatorname{in}})$. Our aim is to provide an analytic form of the Choi matrix of the channel $\mathcal{G}^{\Phi}_{U}$. Utilizing definition of the Choi matrix we obtain:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\left(\mathcal{M}^{\Phi}_{U}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
$=$ $\left(\int_{\operatorname{U}(d)}\operatorname{d}\!U(\mathcal{T}_{U}^{\operatorname{out}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\mathcal{T}_{U^{-1}}^{\operatorname{in}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\right)\rho_{\operatorname{ME}}$
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left(\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}\left(\pi_{\operatorname{U}}^{\operatorname{in}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
Now we can extract the inner unitary representation using the following trick, which holds for arbitrary matrix $X\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$:
$(X^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}(X\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})=({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{X})\rho_{\operatorname{ME}}({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes X^{T}),$ (4)
in which bar denotes complex conjugation. Applying this property we change the order of the inner unitary representations in the tensor product:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\rho_{\operatorname{ME}}\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{T}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
At this point we need the following lemma:
Lemma 1.
Let $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ and $\Psi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ be two channels. Then the following commutation relation holds:
$(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\Psi)=({\rm 1\kern-2.20001ptl}_{\operatorname{out}}\otimes\Psi)\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}).$ (6)
Utilizing the above lemma we obtain:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
The above integral can be always formally expressed in a closed form using generalized duality for group representations. Indeed, note that:
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}},$ (8)
is a finite dimensional regular representation of the unitary group on a representation space $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$. This is because: (i) the contragradient representation [7] $\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)={\pi}_{\operatorname{U}}^{\operatorname{in}}(U^{-1})^{T}$ is a regular representation ([8], Sec. 1.5.2), and (ii) tensor product of two regular representations is a regular representation ([8], ibid). Therefore one can apply the generalized duality theorem for group representations ([8], Sec. 4.2), which boils down to a conclusion, that the action of $\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}$, where $\pi_{\mathcal{C}}$ is a commutant of $\pi_{\operatorname{U}}$ in the algebra of endomorphisms of $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, decomposes uniquely into multiplicity-free direct sum of tensor products of irreducible representations $\xi_{\operatorname{U}}$ of the unitary group $\operatorname{U}(d)$ and irreducible representations $\xi_{\mathcal{C}}$ of a group which is represented by the commutant $\pi_{\mathcal{C}}$:
$\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}=\bigoplus_{k\in\mathcal{I}}\xi_{\operatorname{U}}^{(k)}\otimes\xi_{\mathcal{C}}^{(k)}.$ (9)
[…]
$|.\hat{\Pi}_{kk^{\prime}}^{m_{1}\lambda_{1}m_{2}\lambda_{2}}=\ket{k,m_{1},\lambda_{1}}\!\bra{k^{\prime},m_{2},\lambda_{2}}$. (10)
With the help of the above operators we can define two Schur operator bases:
$\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}$ $=$ $\sum_{m=1}^{D^{k}_{U}}\hat{\Pi}_{kk}^{m\lambda_{1}m\lambda_{2}},$
$\hat{\Pi}^{m_{1}m_{2}}_{k}$ $=$ $\sum_{\lambda=1}^{D^{k}_{C}}\hat{\Pi}_{kk}^{m_{1}\lambda m_{2}\lambda},$ (11)
in which $D^{k}_{U}$ and $D^{k}_{C}$ denote ranges of indices respectively $m$ and $\lambda$. The operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}$ span operator representation of irreducible representations of $\pi_{\mathcal{C}}$. From now on we will be using the notation $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$ in order to explicitly indicate the representation from which the operators are derived. Now, note that the pair $\{\pi_{\operatorname{U}},\pi_{\mathcal{C}}\}$ is a dual reductive pair [8], and as shown in [5] in Theorem 2 one can represent the integral (2) as a projection of the input state (here $J_{\Phi}$) to the basis $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}$ (12)
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\pi_{\operatorname{U}}(U)J_{\Phi}\left(\pi_{\operatorname{U}}(U)\right)^{\dagger}$
$=$ $\sum_{k\in\mathcal{I}}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}].$
In order to utilize this solution one has to explicitly construct operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$ which span irreducible representations of the commutant of $\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}$. Note that the typical choice of the representations on input and output ports of the channel is the collective action of the unitaries on elementary subsystems:
$\pi_{\operatorname{U}}^{\operatorname{in}}(U)$ $=$ $U^{\otimes t_{\operatorname{in}}},$
$\pi_{\operatorname{U}}^{\operatorname{out}}(U)$ $=$ $U^{\otimes t_{\operatorname{out}}},$ (13)
and hence one gets:
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}=U^{\otimes t_{\operatorname{in}}}\otimes\bar{U}^{\otimes t_{\operatorname{out}}}.$ (14)
The commutant of this representation is known to be the so-called walled Brauer algebra [7, 13, 14], the representations of which are difficult to construct [14]. Although there exist implementations of Quantum Schur Transform for this case (so called Mixed Quantum Schur Transform [15]), one can wish to avoid these representations and utilize simpler operator representations of permutations. In order to do so, let us reexpress the averaging operations in a way which removes the complex conjugation. We start from the following lemma:
Lemma 2.
Let $X,Y$ be arbitrary matrices acting on $\mathbb{C}^{r}$ and let $\sigma$ be any state on $\mathbb{C}^{r}\otimes\mathbb{C}^{r}$. Then the following holds:
$(X\otimes\bar{Y})\sigma(X\otimes\bar{Y})^{\dagger}=\left[(X\otimes{Y})\sigma^{\Gamma}(X\otimes{Y})^{\dagger}\right]^{\Gamma},$ (15)
where Γ denotes partial transpose on the second factor.
Proof.
By the linearity argument it suffices to prove the above property for a product state $\sigma=\xi\otimes\tau$. Then we have:
$\left[(X\otimes{Y})(\xi\otimes\tau)^{\Gamma}(X\otimes{Y})^{\dagger}\right]^{\Gamma}$ $=$ $\left[(X\otimes{Y})(\xi\otimes\tau^{T})(X^{\dagger}\otimes{Y}^{\dagger})\right]^{\Gamma}$ (16)
$=$ $\left[(X\xi{X^{\dagger}})\otimes(Y\tau^{T}{Y}^{\dagger})\right]^{\Gamma}$
$=$ $(X\xi{X^{\dagger}})\otimes(Y\tau^{T}{Y}^{\dagger})^{T}$
$=$ $(X\xi{X^{\dagger}})\otimes(\bar{Y}\tau{Y}^{T})$
$=$ $(X\otimes\bar{Y})(\xi\otimes\tau)(X^{\dagger}\otimes{Y}^{T})$
$=$ $(X\otimes\bar{Y})(\xi\otimes\tau)(X\otimes\bar{Y})^{\dagger}.$
Utilizing the above lemma for $X=\pi_{\operatorname{U}}^{\operatorname{out}}(U)$, $Y=\pi_{\operatorname{U}}^{\operatorname{in}}(U)$ and $\sigma=J_{\Phi}$ we obtain:
$J_{\mathcal{M}^{\Phi}_{U}}=\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)J^{\Gamma}_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}\right]^{\Gamma}.$ (17)
For the special case of collective unitaries (2) we have:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)^{\dagger}\right]^{\Gamma}$
$=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
Applying the general formula (12) we have:
$J_{\mathcal{M}^{\Phi}_{U}}$ $=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
$=$ $\left[\mathcal{M}_{U}\left(J_{\Phi}^{\Gamma}\right)\right]^{\Gamma}=\sum_{k}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)^{\Gamma},$
where now the operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]$ are operator representations of permutation operators, and can be constructed directly using the methodology of Young diagrams, see e.g. Appendix of [5] and [16] for a theoretical exposition, and [17, 9, 10, 11] for algorithmic implementations.
3 Averaging quantum channels over non-unitary operations defined via Cartan decomposition
Let us now try to generalize averaging of quantum channels from the unitary averaging to averaging over non-unitary operations described by a non-compact symmetry group $\mathcal{G}$. Such an averaging can be defined via iterated integral over the Cartan decomposition of $\mathcal{G}$, which boils down to a representation of an arbitrary element of $\mathcal{G}$ as a composition of operations from the maximally compact component $\mathcal{K}$ of the group intertwined by an element from a maximal Abelian subgroup $\mathcal{A}$: $G=KAK^{\prime},\,K,K^{\prime}\in\mathcal{K},\,A\in\mathcal{A}$. Such idea of non-compact averaging has been introduced in [18, 5] in the context of averaging quantum states over SLOCC-type operations, and generalized to averaging over arbitrary finite-dimensional matrix representations of reductive groups in the recent work [6]. Following [6], let us take arbitrary matrix representation $\pi_{\mathcal{G}}$ of the group $\mathcal{G}$ and define the averaging of a quantum state with respect to the chosen representation as follows:
$\mathcal{M}_{\mathcal{G}}(\rho)=\int_{\mathcal{K}\times\mathcal{A}\times\mathcal{K}}\left(\pi_{\mathcal{G}}(K)\pi_{\mathcal{G}}(A_{\textrm{n}})\pi_{\mathcal{G}}(K^{\prime})\right)\rho\left(\pi_{\mathcal{G}}(K)\pi_{\mathcal{G}}(A_{\textrm{n}})\pi_{\mathcal{G}}(K^{\prime})\right)^{\dagger}\operatorname{d}\!K\operatorname{d}\!A\operatorname{d}\!K^{\prime}.$ (20)
in which the integration is defined over the Cartan “KAK” decomposition of $\mathcal{G}$. The integration over the two copies of maximally compact components $\mathcal{K},\mathcal{K}^{\prime}$ is performed uniformly, which means that $\int_{\mathcal{K}}\operatorname{d}\!K$ represents Haar integral over $\mathcal{K}$. Integration over $\mathcal{A}$ is performed with respect to any finite normalized measure on this (in general non-compact) manifold, and there appears one crucial assumption: elements of $\mathcal{A}$ enter the integral as normalized matrices: $AA_{\textrm{n}}$=A/||A||, in order to guarantee trace nonincreasing character of the map (20).
Our aim is to define twirling of quantum channels with respect to the representation $\pi_{\mathcal{G}}$ applied as pre- and post-processing operation:
$\mathcal{T}_{G}^{\operatorname{in}}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)^{\dagger},\,\,\sigma_{\operatorname{in}}\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$
$\mathcal{T}_{G}^{\operatorname{out}}(\sigma_{\operatorname{out}})$ $=$ $\pi^{\operatorname{out}}_{{\mathcal{G}}}(G)\sigma_{\operatorname{out}}\pi^{\operatorname{out}}_{{\mathcal{G}}}(G)^{\dagger},\,\,\sigma_{\operatorname{out}}\in\mathcal{B}(\mathcal{H}_{\operatorname{out}}).$ (21)
In order to follow the analogy with (2) we have to define the counterpart of the $\mathcal{T}_{U^{-1}}$ operation from definition (2). Formally every element $G$ from the group $\mathcal{G}$ is invertible, however, the demand of trace-nonincreasing character of a quantum channel forces us to integrate over normalized elements from the group $\mathcal{G}$ in the twirling channel (20). Now, in general normalized matrices, which represent trace-nonincreasing quantum channels, are not invertible as quantum operations. Therefore we define twirling of quantum channels with respect to $\mathcal{G}$ operations as follows:
$\mathcal{M}^{\Phi}_{G}(\rho)=\int\operatorname{d}\!G\,\mathcal{T}_{G}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{G}^{\operatorname{in}\dagger}(\rho),$ (22)
in which we define:
$\mathcal{T}_{G}^{\operatorname{in}\dagger}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{{\mathcal{G}}}(G).$ (23)
Note that the integration in (22) is understood in the sense of Cartan integral (20). In the case of a compact group $\mathcal{G}$ with trivial $\mathcal{A}$ this definition is fully equivalent to (2).
Now all of the derivations (2)–(2) concerning the unitary case hold, and we have the following expression for the Choi matrix of the channel (22):
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int\operatorname{d}\!G\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)\right)J_{\Phi}\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)\right)^{\dagger}.$ (24)
Due to distributivity of the tensor product with respect to ordinary matrix product, we have the following Cartan decompositon:
$\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)=\left(\pi_{\mathcal{G}}^{\operatorname{out}}(K)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(K)\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(A_{\textrm{n}})\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(A_{\textrm{n}})\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(K^{\prime})\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(K^{\prime})\right),$ (25)
in which we omit bar in the middle term since $A_{\textrm{n}}$ is a real diagonal matrix. Since Cartan decomposition commutes with taking representations of the components [6], it can be seen that $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}\equiv\pi_{\mathcal{K}}$, when acting solely on compact elements $K\in\mathcal{K}$ of the group $\mathcal{G}$, is itself a representation of the maximal compact subgroup $\mathcal{K}$ of $\mathcal{G}$. Analogously, in the same sense $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{A}}\equiv\pi_{\mathcal{A}}$, when acting solely on non-compact elements $A\in\mathcal{A}$ can be treated as a representation of the maximal Abelian subgroup $\mathcal{A}$ of $\mathcal{G}$. Hence the integral (24) becomes:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int_{\mathcal{K}\times\mathcal{A}\times\mathcal{K}}\left(\pi_{\mathcal{K}}(K)\pi_{\mathcal{A}}(A_{\textrm{n}})\pi_{\mathcal{K}}(K^{\prime})\right)J_{\Phi}\left(\pi_{\mathcal{K}}(K)\pi_{\mathcal{A}}(A_{\textrm{n}})\pi_{\mathcal{K}}(K^{\prime})\right)^{\dagger}\operatorname{d}\!K\operatorname{d}\!A\operatorname{d}\!K^{\prime}.$ (26)
As shown in [5] such integral decomposes into a product of compact twirling of the input operator and of the square of the non-compact part:
$\mathcal{M}_{\mathcal{G}}(\rho)=\mathcal{M}_{\mathcal{K}}(\rho)\mathcal{M}_{\mathcal{K}}\left(\pi_{\mathcal{A}}(A_{\textrm{n}})^{2}\right),$ (27)
which can be directly represented as:
$\mathcal{M}_{\mathcal{G}}(\rho)=\sum_{k}\left(\frac{\beta_{k}}{D^{k}}\right)\mathcal{M}_{\mathcal{K}}^{(k)}(\rho)=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(\rho\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}],$ (28)
where the coefficients $\beta_{k}$ are integrals over the non-compact components:
[…]
and the operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}]$ are Schur operator basis found by applying generalized duality theorem to the representation $\pi_{\mathcal{K}}=\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$. Therefore, we finally have:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}].$ (30)
In full analogy with the previous case one may want to avoid using contragradient representations in the generalized Schur-Weyl duality. Utilizing Lemma 2 the formula (24) becomes:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int\operatorname{d}\!G\left[\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(G)\right)J_{\Phi}^{\Gamma}\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(G)\right)^{\dagger}\right]^{\Gamma},$ (31)
which can be analogously, using Cartan decomposition, decomposed into:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]\right)^{\Gamma},$ (32)
where now the operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]$ are elements of the Schur operator basis related with decomposing the representations of $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}=\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}$, which are typically much easier to construct. Indeed, let us take as an example the case of $\mathcal{G}=\textrm{SL}(d,\mathbb{C})$, for which the maximally compact Cartan component $\mathcal{K}$ is the special unitary group $\textrm{SU}(d)$. In full analogy to the unitary case we assume collective action of group elements on input and output spaces:
$\pi_{\textrm{SL}}^{\operatorname{in}}(L)$ $=$ $L^{\otimes t_{\operatorname{in}}},\,\,L\in\textrm{SL}(d,\mathbb{C}),$
$\pi_{\textrm{SL}}^{\operatorname{out}}(L)$ $=$ $L^{\otimes t_{\operatorname{out}}},\,\,L\in\textrm{SL}(d,\mathbb{C}).$ (33)
In this case the Choi matrix of a twirled channel reads after (31):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}$ $=$ $\int_{\textrm{SL}(d,\mathbb{C})}\operatorname{d}\!L\left[\left(L^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(L^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}.$ (34)
Now the restricted representation $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$ reads in this case simply $U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}$, for $U\in\textrm{SU}(d)$, and we have, following (32):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)^{\Gamma},$ (35)
where now $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{(t_{\operatorname{in}}+t_{\operatorname{out}})}]$ are ordinary permutation operators corresponding to standard Schur-Weyl duality for $\textrm{U}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$ and the permutation group $\textrm{S}_{(t_{\operatorname{in}}+t_{\operatorname{out}})}$.
4 Dual averaging of channels
In this section we utilize the concept of generalized finite unitary averaging introduced in a recent work [6] in the context of averaging of quantum channels. The idea of generalized finite unitary averaging, originally introduced for averaging quantum states over general symmetry operations, also bases on generalized duality for group representations, however, in analogy to the concept of unitary $t$-designs, it keeps the averaging procedure on the original group representation side, and not on its commutant, as in formulas (12) and (30), in which the Choi matrix of the twirled channel is expressed in terms of operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}$, which span the commutant of tensor product of the original symmetry group. Indeed, the idea in [6] is to represent the Cartan-decomposition-based average of a quantum state with respect to arbitrary symmetry group (20) as a probabilistic mixture of unitary-$1$-designs on invariant subspaces:
$\mathcal{M}_{\mathcal{G}}(\rho)=\sum_{k\in\mathcal{I}}p_{k}\mathcal{D}_{\mathcal{K}}^{(k)}(\rho),$ (36)
in which $\mathcal{D}_{\mathcal{K}}^{(k)}$ represents unitary-$1$-design on $D^{k}_{G}$-dimensional subspace irreducible under the action of the restricted representation $\pi_{\mathcal{K}}$, for $\mathcal{K}$ being the maximal compact subgroup of $\mathcal{G}$:
$\mathcal{D}_{\mathcal{K}}^{(k)}(\rho)=\frac{1}{\left(D^{k}_{G}\right)^{2}}\sum_{l=1}^{\left(D^{k}_{G}\right)^{2}}\tilde{\gamma}_{l}^{(k)}\rho\,\tilde{\gamma}_{l}^{(k)\dagger},$ (37)
and the probabilities $p_{k}$ are specified in full analogy to (28) as $p_{k}=\frac{\beta_{k}}{D^{k}}$. Unitary-$1$-designs are specified by choosing a unitary operator basis $\{\gamma^{(k)}_{l}\}_{l=1}^{\left(D^{k}_{G}\right)^{2}}$ on space $\mathbb{C}^{D^{k}_{G}}$, and further they are extended to act on the entire representation space using basis operators $\hat{\Pi}_{k}^{m_{1}m_{2}}[\pi_{\mathcal{K}}]$ (2) specified by generalized group duality:
$\tilde{\gamma}^{(k)}_{l}=\sum_{m_{1},m_{2}=1}^{D^{k}_{G}}{\left(\gamma^{(k)}_{l}\right)}_{m_{1}m_{2}}\hat{\Pi}^{m_{1}m_{2}}_{k}[\pi_{\mathcal{K}}].$ (38)
Utilizing the above formalism one can express the Choi matrix $J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ of a twirled channel (24) in a design-like form:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k\in\mathcal{I}}\left(\frac{\beta_{k}}{D^{k}}\right)\mathcal{D}_{\mathcal{K}}^{(k)}(J_{\Phi})=\sum_{k\in\mathcal{I}}\left(\frac{\beta_{k}}{D^{k}}\right)\frac{1}{\left(D^{k}_{G}\right)^{2}}\sum_{l=1}^{\left(D^{k}_{G}\right)^{2}}\tilde{\gamma}_{l}^{(k)}J_{\Phi}\,\tilde{\gamma}_{l}^{(k)\dagger},$ (39)
where the operators (38) are defined for $\hat{\Pi}^{m_{1}m_{2}}_{k}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{K}}^{\operatorname{in}}]$. In principle in the above formulas one can use arbitrary unitary operator basis, however the typical choice due to its simplicity of construction is the Heisenberg-Weyl unitary operator bases [19], comprising operators:
$\{\omega_{D^{k}_{G}}^{ij}Z_{D^{k}_{G}}^{i}X_{D^{k}_{G}}^{j}\}_{i,j=0}^{{D^{k}_{G}}-1}$
in which the complex root of unity reads: $\omega_{{D^{k}_{G}}}=e^{2i\pi/{D^{k}_{G}}}$, and the operators $Z,X$ are unitary generalizations of Pauli matrices $\sigma_{z},\sigma_{x}$.
5 Channel $t$-designs
The methods of averaging quantum channels described in previous sections are based on generalized group duality and therefore, though finite, they demand performing projections on irreducible subspaces related with representations of either the original symmetry group for input and output ports or the commutant of its action. One may expect a finite averaging method which respects the operator structure of the original group actions, as in the case of unitary $t$-designs. We will provide such a generalization utilizing Lemma 2 for the following case:
$\pi_{\mathcal{G}}^{\operatorname{in}}(G)$ $=$ $G^{\otimes t_{\operatorname{in}}},\,\,G\in\mathcal{G},$
$\pi_{\mathcal{G}}^{\operatorname{out}}(G)$ $=$ $G^{\otimes t_{\operatorname{out}}},\,\,G\in\mathcal{G},$ (40)
in which $\mathcal{G}$ is an arbitrary subgroup of $\textrm{GL}(d,\mathbb{C})$. We utilize the following lemma:
Lemma 3.
Let $\{G_{i},w_{i}\}$ be a (weighted) group $t$-design, namely:
$\int_{\mathcal{G}}\operatorname{d}\!G\,G^{\otimes t}\rho\,G^{\otimes t\dagger}=\sum_{i\in\mathcal{X}}w_{i}G_{i}^{\otimes t}\rho\,G_{i}^{\otimes t\dagger}.$ (41)
Then the following holds:
$\int_{\mathcal{G}}\operatorname{d}\!G\,\left[G^{\otimes t}\rho^{\Gamma}\,G^{\otimes t\dagger}\right]^{\Gamma}=\sum_{i\in\mathcal{X}}w_{i}\left[G_{i}^{\otimes t}\rho^{\Gamma}\,G_{i}^{\otimes t\dagger}\right]^{\Gamma}.$ (42)
Proof.
The above holds due to: (i) linearity of partial transpose, (ii) the fact that (41) holds for any $\rho$ (in fact $\rho$ is redundant in this formula). ∎
Utilizing the above lemma we can define channel $t$-design, where $t=t_{\operatorname{in}}+t_{\operatorname{out}}$ as a finite set of matrices and weights $\{G_{i},w_{i}\}$ which reconstruct the Choi matrix of an averaged channel:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int_{\mathcal{G}}\operatorname{d}\!G\left[\left(G^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(G^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}$ (43)
$=$ $\sum_{i\in\mathcal{X}}w_{i}\left[\left(G_{i}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(G_{i}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}.$
The above construction is very general and can be applied whenever one can construct $t$-design for a given symmetry group $\mathcal{G}$ and demanded value of $t=t_{\operatorname{in}}+t_{\operatorname{out}}$. In the case of $\mathcal{G}$ being a unitary group there is a huge literature on this topic [20, 21, 22, 23, 24, 25], whereas in the case of $\mathcal{G}=\textrm{SL}(2,\mathbb{C})$ one can construct respective designs as product designs comprising unitary ones and Gauss-type quadratues, see [18] for detailed presentation.
Relation to designs for channels.
Several notions of designs for channels have been proposed in the literature, often by pushing forward a design on a larger unitary group through a fixed Stinespring dilation or by imposing moment-matching conditions on ensembles of CPTP maps, see e.g. [26]. The channel $t$-design notion in Eq. (43) is tailored to the present twirling problem: it is the minimal finite condition required to reproduce the averaged Choi operator for collective pre/post actions, and it follows directly from the partial-transpose reduction and any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$. This perspective complements other channel-design frameworks by making the symmetry and the induced representation on the Choi space explicit.
6 Conclusions and outlook
We developed a representation-theoretic framework for averaging (twirling) of quantum channels that is driven by channel–state duality. Starting from the definition of channel twirling as uniform pre- and post-processing by the symmetry group representations on the input and output spaces, we derived an explicit Choi-level formulation in which the twirl becomes a group average acting directly on the Choi operator under the induced representation $\pi^{\mathrm{out}}\otimes\bar{\pi}^{\mathrm{in}}$. This viewpoint identifies the twirl as a projection onto the commutant algebra and leads to closed-form expressions in terms of generalized Schur operator bases obtained from generalized Schur-Weyl duality.
A central simplification arises in the collective setting $\pi^{\mathrm{in}}(U)=U^{\otimes t_{\mathrm{in}}}$ and $\pi^{\mathrm{out}}(U)=U^{\otimes t_{\mathrm{out}}}$, where the commutant of $U^{\otimes t_{\mathrm{out}}}\otimes\bar{U}^{\otimes t_{\mathrm{in}}}$ is the walled Brauer algebra. By exploiting a partial-transpose identity, we showed that channel twirling can be reduced to an ordinary Schur-Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\mathrm{in}}+t_{\mathrm{out}})}$. This removes the contragredient action at the cost of a final partial transpose and enables practical evaluation using permutation-operator technology rather than mixed Schur transforms or explicit walled Brauer idempotents.
We further extended the construction beyond compact unitary symmetries to averaging over non-unitary operations described by reductive groups via Cartan decomposition. The resulting twirl admits a decomposition into invariant-sector projections weighted by coefficients $\beta_{k}$ determined solely by the Abelian Cartan component, and an analogous partial-transpose reduction applies to avoid contragredient representations on the compact part. In addition, we discussed a dual averaging realization in which the (Cartan) twirl can be implemented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors through embedded unitary operator bases. Finally, motivated by unitary designs, we introduced a notion of channel $t$-design for collective actions: any (weighted) group $t$-design for $t=t_{\mathrm{in}}+t_{\mathrm{out}}$ induces a finite reconstruction of the averaged Choi operator, in this case preserving the structure of the original symmetry group action.
Several directions follow naturally. On the constructive side, it is important to develop efficient exact and approximate finite averaging sets for broader classes of symmetry groups and representations, and to quantify the error incurred when replacing Haar/Cartan integrals by designs in the induced channel-$t$ setting. On the algorithmic side, the partial-transpose reduction suggests scalable implementations of channel twirls based on permutation-operator decompositions and Schur-Weyl projectors, with complexity controlled by $t_{\mathrm{in}}+t_{\mathrm{out}}$ rather than the full walled Brauer machinery. Finally, extending the present approach to approximate symmetries, to non-identical local dimensions, and to other physically motivated group actions (beyond collective tensor powers) may further broaden the range of channel-averaging problems that admit analytic reduction and finite, design-like realizations.
Acknowledgements
MM acknowledges support from the National Science Center (NCN), Poland, under Project Opus No. 2024/53/B/ST2/02026. ZP and ŁP acknowledge support from the National Science Center (NCN), Poland, under Project Opus No. 2022/47/B/ST6/02380.