### v2

# Choi-level twirling of quantum channels: finite constructions and non-compact transformations

Abstract
Twirling, i.e. averaging over symmetry actions, is a standard tool for reducing quantum states and channels to a symmetry-invariant form. We study channel twirling from the perspective of the channel–state duality and provide a constructive Choi-level description of the averaging map induced by arbitrary input/output representations. Our main technical result concerns the collective setting: for $\pi^{\operatorname{in}}(U)=U^{\otimes t_{\operatorname{in}}}$ and $\pi^{\operatorname{out}}(U)=U^{\otimes t_{\operatorname{out}}}$, we introduce a partial-transpose reduction that removes the contragredient action and converts the mixed (walled Brauer) channel twirl into an ordinary Schur–Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, enabling explicit permutation-based formulas without constructing walled Brauer idempotents or mixed Schur transforms. Beyond compact symmetries, we extend channel twirling to reductive, generally non-unitary groups via Cartan decomposition and obtain an invariant-sector decomposition of the averaged Choi operator with weights determined solely by the Abelian Cartan component. Finally, we provide two finite realizations of channel averaging: a “dual” implementation as a convex mixture of unitary-$1$-design channels acting on invariant sectors, and a design-like reconstruction showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\operatorname{in}}+t_{\operatorname{out}}$.
1 Introduction
1.1 General Motivation
Averaging procedures based on symmetry groups play a central role in quantum information theory [1, 2]. For quantum states, uniform averaging (twirling) with respect to a group action provides a systematic way of reducing descriptions to symmetry-invariant data, with applications ranging from entanglement theory to randomized benchmarking and noise-invariant encoding [3, 4, 5, 6, 7, 8]. For quantum channels, averaging is typically defined operationally via uniform pre- and post-processing by symmetry transformations, leading to effective noise models such as depolarizing or stochastic channels [2, 9, 10, 11, 12, 13].
Despite its widespread use, channel averaging is most often treated at the level of operational definitions and specific examples. The Choi–Jamiołkowski isomorphism provides a natural representation of quantum channels as bipartite operators. While this framework is invaluable for tasks such as defining operational distance measures [14] and discriminating quantum operations [15], a general constructive Choi-level formula that works uniformly for arbitrary collective tensor actions and input/output transformations remains largely absent from the literature. Consequently, practical evaluation often relies on case-by-case representation-theoretic machinery. Existing approaches either focus on special classes of symmetries, such as finite or compact groups with irreducible representations, or rely on case-by-case constructions tailored to specific physical settings.
A natural expectation is that channel twirling should correspond to a projection onto symmetry-invariant subspaces of the Choi operator. However, making this correspondence precise for arbitrary input and output transformations, and doing so in a way that is both conceptually transparent and practically usable, turns out to be nontrivial. This difficulty becomes particularly apparent in collective scenarios, where channels act on multiple identical subsystems and the relevant symmetry algebra is no longer given by the symmetric group alone (as in the case of twirling of quantum states) [8], but by the commutant of the tensor product action $U^{\otimes p}\otimes\bar{U}^{\otimes q}$, namely the walled Brauer algebra [16].
1.2 State of the art: covariant and equivariant channels, and mixed Schur–Weyl tools.
Symmetry constraints on quantum channels are most commonly formulated in terms of covariance [^1] It is important to distinguish two related but different notions: covariance/equivariance is a symmetry constraint defining a class of channels, whereas twirling is an averaging procedure that maps an arbitrary channel to a symmetric one.. A channel $\Phi:\mathcal{B}(\mathcal{H}_{\mathrm{in}})\to\mathcal{B}(\mathcal{H}_{\mathrm{out}})$ is called $\mathcal{G}$-covariant if there exist (typically unitary) representations $\pi_{\mathcal{G}}^{\textrm{in}}$ and $\pi_{\mathcal{G}}^{\textrm{out}}$ of the symmetry group $\mathcal{G}$, such that
$\Phi\bigl(\pi_{\mathcal{G}}^{\textrm{in}}(g)\,\rho\,\pi_{\mathcal{G}}^{\textrm{in}}(g)^{\dagger}\bigr)=\pi_{\mathcal{G}}^{\textrm{out}}(g)\,\Phi(\rho)\,\pi_{\mathcal{G}}^{\textrm{out}}(g)^{\dagger}\qquad\forall g\in\mathcal{G}.$ (1)
It is a well-known consequence of the Choi–Jamiołkowski isomorphism that this covariance condition is equivalent to a commutation relation at the Choi level [17, 10, 18]:
$\bigl[\,\pi_{\mathcal{G}}^{\textrm{out}}(g)\otimes\bar{\pi}_{\mathcal{G}}^{\textrm{in}}(g),\,J_{\Phi}\,\bigr]=0\qquad\forall g\in\mathcal{G},$ (2)
that is, the Choi operator must lie in the commutant of the induced representation $\pi_{\mathcal{G}}^{\textrm{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\textrm{in}}$.
When the input and output representations are identical, or related in a direct “same-action” manner (e.g. $U^{\otimes p}$ on the input and $U^{\otimes q}$ on the output for $U\in\mathrm{U}(d)$), the literature often uses the term equivariance to emphasize that applying the symmetry to the input corresponds to applying (essentially) the same symmetry to the output. This distinction is especially natural in the collective setting of many identical subsystems, where one studies channels whose Choi operators commute with the tensor action $U^{\otimes q}\otimes\bar{U}^{\otimes p}$.
A substantial body of work analyzes the structure and applications of such symmetry-restricted channels. For finite groups, a detailed characterization is available in the irreducibly covariant case, where representation-theoretic data yields explicit decompositions of the Choi operator and necessary and sufficient conditions for complete positivity and trace preservation [17, 19]. For continuous unitary symmetries and multipartite settings, mixed Schur–Weyl duality identifies the commutant of $U^{\otimes q}\otimes\bar{U}^{\otimes p}$ with the walled Brauer algebra, providing a powerful language for block diagonalization and parametrization [20, 18, 21]. This perspective underlies modern optimization frameworks: unitary-equivariant semidefinite programs can be reduced to linear programs by expressing feasible operators in a walled Brauer basis and exploiting additional symmetry constraints [22, 18]. On the quantum-algorithmic side, efficient circuit constructions for the mixed Schur transform enable implementations and manipulations of unitary-equivariant channels in a variety of operational settings [20, 23].
In parallel, channel twirling is a standard operational tool, defined by averaging pre- and post-processing operations over a group action [2, 9, 10, 11, 12, 13]. Connections to unitary designs are well understood in several important cases. In particular, twirling by a unitary $1$-design yields a stochastic channel, although the resulting channel may depend on the specific choice of the design [11]. Much of the literature, however, focuses on concrete symmetry models or assumes the invariant decomposition a priori, rather than providing a general and constructive description of channel twirling directly at the level of the Choi operator.
Despite extensive work on covariant channels and mixed Schur–Weyl duality [20, 22, 24, 18, 21, 13], a general and constructive description of channel twirling at the Choi level, especially beyond compact symmetries, has been missing. Moreover in much of the literature on this topic the symmetry is imposed as a structural constraint on $\Phi$ (equivariant/covariant channels), whereas here we focus on the averaging map that sends an arbitrary $\Phi$ to its symmetry-averaged counterpart and we make this map constructive at the Choi level.
1.3 Outline
In this work, we develop a unified framework for averaging quantum channels based on the channel–state duality. Starting from the standard definition of channel twirling as uniform pre- and post-processing by symmetry operations, we show that channel averaging can be expressed exactly as a group twirl acting directly on the Choi operator. More precisely, the Choi operator of a twirled channel is obtained by averaging the original Choi operator with respect to the induced representation on the joint output–input space.
Although this correspondence between operational twirling and a Choi-level group action appears implicitly or explicitly in previous works, we elevate it to a central structural principle. In our approach, the Choi-level formulation is not merely a reformulation, but the fundamental tool that enables a uniform treatment of arbitrary representations of input and output symmetry transformations, collective tensor actions, and non-compact operations.
Viewed in this way, channel twirling becomes an explicit projection onto the commutant of the induced symmetry representation, placing channel averaging exactly within the framework of generalized Schur–Weyl duality [25, 26, 27, 6]. This structural perspective is what allows us to identify invariant-sector decompositions, derive finite-sum formulas, and extend the formalism beyond the compact unitary setting.
In the important collective setting, where the symmetry group acts identically on multiple input and output subsystems, the commutant algebra is given by the walled Brauer algebra. To avoid the technical complexity associated with constructing explicit Brauer idempotents, we introduce a partial-transpose reduction that maps channel twirling to an ordinary Schur–Weyl twirl of a partially transposed Choi operator. As a result, the evaluation of collective channel twirls can be reduced to standard permutation-based formulas, enabling explicit formulas in terms of permutation operators and Schur projectors, without constructing walled Brauer idempotents or mixed Schur transforms.
Genuinely beyond the symmetries represented by compact groups, we extend our framework to averaging over arbitrary finite-dimensional representations of reductive, generally non-compact groups. Using Cartan decomposition, we show that the resulting channel twirl admits a decomposition into invariant-sector projections with weights determined solely by the Abelian component of the group. This provides a controlled and well-defined notion of channel averaging beyond the unitary setting, where naive uniform averaging is not available.
Finally, we address finite realizations of channel averaging. We introduce a dual averaging representation in which the twirled channel is expressed as a convex mixture of unitary $1$-design channels acting on invariant sectors. We further establish a design-like reconstruction result showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\mathrm{in}}+t_{\mathrm{out}}$, allowing exact reconstruction of the twirled Choi operator from a finite sum.
The results presented here provide a general and constructive description of channel averaging that unifies and extends existing approaches. They offer both conceptual clarification of the role of symmetry in quantum channels and practical tools for analyzing and implementing averaged channels in a wide range of settings.
1.4 Notation
Firstly let us introduce notation inspired by previous works on averaging of quantum states [8, 28]. In this work we assume that the dimensions of elementary systems constituting input and output subspaces are equal, however the number of elementary subsystems can be different for input and output spaces. Hence the input and output Hilbert spaces are specifed respectively as: $\mathcal{H}_{\operatorname{in}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{in}}}$ and $\mathcal{H}_{\operatorname{out}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{out}}}$. We consider arbitrary channels $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ mapping $t_{\operatorname{in}}$ qu-$d$-its to $t_{\operatorname{out}}$ qu-$d$-its.
Following [28] let us introduce two arbitrary unitary representations $\pi_{\operatorname{U}}^{\operatorname{in}}$ and $\pi_{\operatorname{U}}^{\operatorname{out}}$ of a unitary group $\operatorname{U}(d)$, acting respectively on $\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ and $\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ via maps:
$\mathcal{T}_{U}^{\operatorname{in}}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{\operatorname{U}}(U)\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger},\,\,\sigma_{\operatorname{in}}\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$
$\mathcal{T}_{U}^{\operatorname{out}}(\sigma_{\operatorname{out}})$ $=$ $\pi^{\operatorname{out}}_{\operatorname{U}}(U)\sigma_{\operatorname{out}}\pi^{\operatorname{out}}_{\operatorname{U}}(U)^{\dagger},\,\,\sigma_{\operatorname{out}}\in\mathcal{B}(\mathcal{H}_{\operatorname{out}}).$ (3)
Clearly we have: $\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\sigma_{\operatorname{in}})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)$, since $\pi^{\operatorname{in}}_{\operatorname{U}}(U^{-1})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{-1}=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}$. The last property holds because any representation is a group homomorphism. Following ideas of [2, 9, 11, 18], let us define twirling of quantum channels as follows:
$\mathcal{M}^{\Phi}_{\operatorname{U}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\left(\Phi\left(\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)\right)\right),$ (4)
in which the integration is performed with respect to a Haar measure on the unitary group. Further on we will generalize the above averaging procedure to the case of non-compact transformations.
1.5 Main results
Our contributions can be summarized as follows.
  1. 1.
    In the collective setting, we introduce a partial-transpose reduction that removes the contragredient action and converts the mixed Schur–Weyl (walled Brauer) twirl into an ordinary Schur–Weyl twirl on $t_{\operatorname{in}}+t_{\operatorname{out}}$ tensor factors, enabling permutation-based formulas without constructing walled Brauer idempotents.
  2. 2.
    We extend channel twirling beyond compact symmetries to reductive (generally non-compact) groups via Cartan decomposition, obtaining a weighted invariant-sector decomposition with weights determined solely by the Abelian Cartan component.
  3. 3.
    We provide two finite realizations of the channel averaging process: a “dual averaging” implementation as a convex mixture of unitary-$1$-design channels on invariant sectors, and a $t$-design-like reconstruction result showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\operatorname{in}}+t_{\operatorname{out}}$.
The structural Choi-level identity underpinning channel twirling (Lemma 1) records the standard induced action arising from pre-/post-processing and channel–state duality. Our main technical novelty begins in the collective setting, where Lemma 3 yields a partial-transpose reduction that removes the contragredient action and converts the mixed (walled Brauer) twirl into an ordinary Schur–Weyl twirl on $t_{\operatorname{in}}+t_{\operatorname{out}}$ tensor factors. This enables explicit permutation-operator formulas without constructing walled Brauer idempotents or invoking mixed Schur transforms.
Partial-transpose reduction (collective setting).
Theorem 1 (Collective channel twirl reduces to an ordinary Schur–Weyl twirl, Section 2).
Assume the collective representations $\pi_{\mathrm{U}}^{\operatorname{in}}(U)=U^{\otimes t_{\operatorname{in}}}$ and $\pi_{\mathrm{U}}^{\operatorname{out}}(U)=U^{\otimes t_{\operatorname{out}}}$. Let $\Gamma$ denote partial transpose on the input factor of $\mathcal{H}_{\operatorname{out}}\otimes\mathcal{H}_{\operatorname{in}}$. Then the Choi operator of the Haar-twirled channel satisfies
$J_{\mathcal{M}^{\Phi}_{\mathrm{U}}}=\left[\int_{\operatorname{U}(d)}\mathrm{d}U\;\Big(U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\Big)\,J_{\Phi}^{\Gamma}\,\Big(U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\Big)^{\dagger}\right]^{\Gamma}.$
Consequently, evaluating $\mathcal{M}^{\Phi}_{\mathrm{U}}$ can be reduced to a standard Schur–Weyl projection of $J_{\Phi}^{\Gamma}$, so one may work directly with permutation operators on $(t_{\operatorname{in}}+t_{\operatorname{out}})$ tensor factors.
Beyond compact symmetries: reductive (Cartan) twirl.
Theorem 2 (Reductive (Cartan) channel twirl and invariant-sector decomposition, Section 3).
Let $\mathcal{G}$ be a reductive group with Cartan decomposition $\mathcal{G}=\mathcal{K}\,\mathcal{A}\,\mathcal{K}$, and let $\pi_{\mathcal{G}}^{\operatorname{in}/\operatorname{out}}$ be finite-dimensional matrix representations on $\mathcal{H}_{\operatorname{in}/\operatorname{out}}$. Define channel twirling by integrating over $\mathcal{K}\times\mathcal{A}\times\mathcal{K}$ (with normalized Abelian component) and using the adjoint action on the input side. Then the averaged Choi operator admits an invariant-sector decomposition
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\mathcal{P}_{k}(J_{\Phi}),$
where $\{\mathcal{P}_{k}\}$ are commutant-sector projections determined by the restriction of $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}$ to the maximal compact subgroup $\mathcal{K}$, while the coefficients $\beta_{k}$ depend only on the Abelian component $\mathcal{A}$ via a real-valued integral. Moreover, in the collective case one may again express $\mathcal{P}_{k}$ using ordinary Schur–Weyl permutation operators after applying the partial transpose trick.
Finite realizations: dual averaging and induced channel designs.
Theorem 3 (Dual averaging as a mixture of unitary-$1$-design channels, Section 4).
The Cartan twirl $\mathcal{M}_{\mathcal{G}}^{\Phi}$ can be represented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors:
$J_{\mathcal{M}_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\frac{1}{(D_{k}^{G})^{2}}\sum_{l=1}^{(D_{k}^{G})^{2}}\tilde{\gamma}^{(k)}_{l}\,J_{\Phi}\,\tilde{\gamma}^{(k)\dagger}_{l}$,
where $\{\tilde{\gamma}^{(k)}_{l}\}$ are Kraus operators obtained by embedding a unitary operator basis on respective irreducible subspaces into the full representation space.
Proposition 1 (Channel $t$-design induced by a weighted group $t$-design, Section 5).
Let $t=t_{\operatorname{in}}+t_{\operatorname{out}}$ and suppose $\{(G_{i},w_{i})\}$ is a weighted group $t$-design for a subgroup $\mathcal{G}\subseteq\mathrm{GL}(d,\mathbb{C})$. Then the averaged Choi operator can be reconstructed from a finite sum:
$J_{\mathcal{M}_{\mathcal{G}}^{\Phi}}=\left[\int_{\mathcal{G}}\mathrm{d}G\;G^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G^{\otimes t})^{\dagger}\right]^{\Gamma}=\left[\sum_{i}w_{i}\;G_{i}^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G_{i}^{\otimes t})^{\dagger}\right]^{\Gamma}.$
Warm-up identity: induced Choi action.
For completeness, we also record the basic Choi-level identity that relates pre/post twirling of channels to the induced action on the Choi operator (Section 2). Since variants of this statement appear in several contexts in the literature, we treat it as a structural lemma underpinning the subsequent constructions.
Lemma 1 (Choi form of channel twirling under an induced representation).
Let $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\to\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ be a quantum channel with Choi operator $J_{\Phi}$. Given unitary representations $\pi_{\mathrm{U}}^{\operatorname{in}}$ on $\mathcal{H}_{\operatorname{in}}$ and $\pi_{\mathrm{U}}^{\operatorname{out}}$ on $\mathcal{H}_{\operatorname{out}}$, define
$\mathcal{M}_{\mathrm{U}}^{\Phi}(\rho)=\int_{\operatorname{U}(d)}dU\;\mathcal{T}^{\operatorname{out}}_{U}\circ\Phi\circ\mathcal{T}^{\operatorname{in}}_{U^{-1}}(\rho),\qquad\mathcal{T}^{\operatorname{in}/\operatorname{out}}_{U}(X)=\pi_{\mathrm{U}}^{\operatorname{in}/\operatorname{out}}(U)\,X\,\pi_{\mathrm{U}}^{\operatorname{in}/\operatorname{out}}(U)^{\dagger}.$
Then
$J_{\mathcal{M}_{\mathrm{U}}^{\Phi}}=\int_{\operatorname{U}(d)}dU\;\Big(\pi_{\mathrm{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\mathrm{U}}^{\operatorname{in}}(U)\Big)\,J_{\Phi}\,\Big(\pi_{\mathrm{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\mathrm{U}}^{\operatorname{in}}(U)\Big)^{\dagger}.$
2 Averaging quantum channels over arbitrary unitary transformations
In this section we provide an analytic formula for the action of the channel (4) utilizing channel-state duality. Let $J_{\Phi}=(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left(\rho_{\operatorname{ME}}\right)\in\mathcal{B}(\mathcal{H}_{\operatorname{out}})\otimes\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ denote the Choi matrix of the channel $\Phi$, in which $\rho_{\operatorname{ME}}$ is a non-normalized density matrix of a maximally entangled state on $\mathcal{H}_{\operatorname{in}}\otimes\mathcal{H}_{\operatorname{in}}$ and ${\rm 1\kern-2.20001ptl}_{\operatorname{in}}$ is the identity channel on $\mathcal{B}(\mathcal{H}_{\operatorname{in}})$. Our aim is to provide an analytic form of the Choi matrix of the channel $\mathcal{G}^{\Phi}_{U}$. Utilizing definition of the Choi matrix we obtain:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\left(\mathcal{M}^{\Phi}_{\operatorname{U}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
$=$ $\left(\int_{\operatorname{U}(d)}\operatorname{d}\!U(\mathcal{T}_{U}^{\operatorname{out}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\mathcal{T}_{U^{-1}}^{\operatorname{in}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\right)\rho_{\operatorname{ME}}$
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left(\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}\left(\pi_{\operatorname{U}}^{\operatorname{in}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
Now we can extract the inner unitary representation using the following trick, which holds for arbitrary matrix $X\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$:
$(X^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}(X\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})=({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{X})\rho_{\operatorname{ME}}({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes X^{T}),$ (6)
in which bar denotes complex conjugation. Applying this property we change the order of the inner unitary representations in the tensor product:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\rho_{\operatorname{ME}}\left({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)^{T}\right)\right]\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
At this point we need the following lemma:
Lemma 2.
Let $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ and $\Psi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{in}})$ be two channels. Then the following commutation relation holds:
$(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ({\rm 1\kern-2.20001ptl}_{\operatorname{in}}\otimes\Psi)=({\rm 1\kern-2.20001ptl}_{\operatorname{out}}\otimes\Psi)\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}).$ (8)
Utilizing the above lemma we obtain:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\rho_{\operatorname{ME}}\right)\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}.$
The above integral can be always formally expressed in a closed form using generalized duality for group representations. Indeed, note that:
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}},$ (10)
is a finite dimensional rational representation of the unitary group on a representation space $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$. This is because: (i) the contragradient representation [17] $\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)={\pi}_{\operatorname{U}}^{\operatorname{in}}(U^{-1})^{T}$ is a rational representation ([29], Sec. 1.5.2), and (ii) tensor product of two rational representations is a rational representation ([29], ibid). Therefore one can apply the generalized duality theorem for group representations ([29], Sec. 4.2.1) [^2] Note that the assumption of local regularity in the general duality theorem is in the finite dimensional case equivalent to rationality of the representation, see Sec. 1.5.1 of [29]., which boils down to a conclusion, that the action of $\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}$, where $\pi_{\mathcal{C}}$ is a commutant of $\pi_{\operatorname{U}}$ in the algebra of endomorphisms of $(\mathbb{C}^{d})^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, decomposes uniquely into multiplicity-free direct sum of tensor products of irreducible representations $\xi_{\operatorname{U}}$ of the unitary group $\operatorname{U}(d)$ and irreducible representations $\xi_{\mathcal{C}}$ of a group which is represented by the commutant $\pi_{\mathcal{C}}$:
$\pi_{\operatorname{U}}\times\pi_{\mathcal{C}}=\bigoplus_{k\in\mathcal{I}}\xi_{\operatorname{U}}^{(k)}\otimes\xi_{\mathcal{C}}^{(k)}.$ (11)
[…]
$|.\hat{\Pi}_{kk^{\prime}}^{m_{1}\lambda_{1}m_{2}\lambda_{2}}=\ket{k,m_{1},\lambda_{1}}\!\bra{k^{\prime},m_{2},\lambda_{2}}$. (12)
With the help of the above operators we can define two Schur operator bases:
$\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}$ $=$ $\sum_{m=1}^{D^{k}_{U}}\hat{\Pi}_{kk}^{m\lambda_{1}m\lambda_{2}},$
$\hat{\Pi}^{m_{1}m_{2}}_{k}$ $=$ $\sum_{\lambda=1}^{D^{k}_{C}}\hat{\Pi}_{kk}^{m_{1}\lambda m_{2}\lambda},$ (13)
in which $D^{k}_{U}$ and $D^{k}_{C}$ denote ranges of indices respectively $m$ and $\lambda$. The operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}$ span operator representation of irreducible representations of $\pi_{\mathcal{C}}$. From now on we will be using the notation $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$ in order to explicitly indicate the representation from which the operators are derived. Now, note that the pair $\{\pi_{\operatorname{U}},\pi_{\mathcal{C}}\}$ is a dual reductive pair [29], and as shown in [8] in Theorem 2 one can represent the integral (2) as a projection of the input state (here $J_{\Phi}$) to the basis $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)J_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}$ (14)
$=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\pi_{\operatorname{U}}(U)J_{\Phi}\left(\pi_{\operatorname{U}}(U)\right)^{\dagger}$
$=$ $\sum_{k\in\mathcal{I}}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}].$
In order to utilize this solution one has to explicitly construct operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}]$ which span irreducible representations of the commutant of $\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}$. Note that the typical choice of the representations on input and output ports of the channel is the collective action of the unitaries on elementary subsystems:
$\pi_{\operatorname{U}}^{\operatorname{in}}(U)$ $=$ $U^{\otimes t_{\operatorname{in}}},$
$\pi_{\operatorname{U}}^{\operatorname{out}}(U)$ $=$ $U^{\otimes t_{\operatorname{out}}},$ (15)
and hence one gets:
$\pi_{\operatorname{U}}=\pi_{\operatorname{U}}^{\operatorname{out}}\otimes\bar{\pi}_{\operatorname{U}}^{\operatorname{in}}=U^{\otimes t_{\operatorname{out}}}\otimes\bar{U}^{\otimes t_{\operatorname{in}}}.$ (16)
The commutant of this representation is known to be the so-called walled Brauer algebra [17, 19, 21], the representations of which are difficult to construct [21]. Although there exist implementations of Quantum Schur Transform for this case (so called Mixed Quantum Schur Transform [20, 22, 24]), one can wish to avoid these representations and utilize simpler operator representations of permutations. In order to do so, let us reexpress the averaging operations in a way which removes the complex conjugation. We start from the following lemma:
Lemma 3.
Let $X,Y$ be arbitrary matrices acting on $\mathbb{C}^{r}$ and let $\sigma$ be any state on $\mathbb{C}^{r}\otimes\mathbb{C}^{r}$. Then the following holds:
$(X\otimes\bar{Y})\sigma(X\otimes\bar{Y})^{\dagger}=\left[(X\otimes{Y})\sigma^{\Gamma}(X\otimes{Y})^{\dagger}\right]^{\Gamma},$ (17)
where Γ denotes partial transpose on the second factor.
Proof.
By the linearity argument it suffices to prove the above property for a product state $\sigma=\xi\otimes\tau$. Then we have:
$\left[(X\otimes{Y})(\xi\otimes\tau)^{\Gamma}(X\otimes{Y})^{\dagger}\right]^{\Gamma}$ $=$ $\left[(X\otimes{Y})(\xi\otimes\tau^{T})(X^{\dagger}\otimes{Y}^{\dagger})\right]^{\Gamma}$ (18)
$=$ $\left[(X\xi{X^{\dagger}})\otimes(Y\tau^{T}{Y}^{\dagger})\right]^{\Gamma}$
$=$ $(X\xi{X^{\dagger}})\otimes(Y\tau^{T}{Y}^{\dagger})^{T}$
$=$ $(X\xi{X^{\dagger}})\otimes(\bar{Y}\tau{Y}^{T})$
$=$ $(X\otimes\bar{Y})(\xi\otimes\tau)(X^{\dagger}\otimes{Y}^{T})$
$=$ $(X\otimes\bar{Y})(\xi\otimes\tau)(X\otimes\bar{Y})^{\dagger}.$
Utilizing the above lemma for $X=\pi_{\operatorname{U}}^{\operatorname{out}}(U)$, $Y=\pi_{\operatorname{U}}^{\operatorname{in}}(U)$ and $\sigma=J_{\Phi}$ we obtain:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}=\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)J^{\Gamma}_{\Phi}\left(\pi_{\operatorname{U}}^{\operatorname{out}}(U)\otimes\pi_{\operatorname{U}}^{\operatorname{in}}(U)\right)^{\dagger}\right]^{\Gamma}.$ (19)
For the special case of collective unitaries (2) we have:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\int_{\operatorname{U}(d)}\operatorname{d}\!U\left[\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes t_{\operatorname{out}}}\otimes{U}^{\otimes t_{\operatorname{in}}}\right)^{\dagger}\right]^{\Gamma}$
$=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
Applying the general formula (14) we have:
$J_{\mathcal{M}^{\Phi}_{\operatorname{U}}}$ $=$ $\left[\int_{\operatorname{U}(d)}\operatorname{d}\!U\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)J_{\Phi}^{\Gamma}\left(U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}\right)^{\dagger}\right]^{\Gamma}.$
$=$ $\sum_{k}\frac{1}{D^{k}_{U}}\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]\right)^{\Gamma},$
where now the operators $\hat{\Lambda}^{\lambda_{1}\lambda_{2}}_{k}[U^{\otimes(t_{\operatorname{out}}+t_{\operatorname{in}})}]$ are operator representations of permutation operators, and can be constructed directly using the methodology of Young diagrams, see e.g. Appendix of [8] and [35] for a theoretical exposition, and [30, 31, 32, 33] for algorithmic implementations.
3 Averaging quantum channels over non-unitary operations defined via Cartan decomposition
Let us now try to generalize averaging of quantum channels from the unitary averaging to averaging over finite-dimensional non-unitary representations of a non-compact symmetry group $\mathcal{G}$. Such an averaging can be defined via iterated integral over the Cartan decomposition of $\mathcal{G}$, which boils down to a representation of an arbitrary element of $\mathcal{G}$ as a composition of operations from the maximally compact component $\mathcal{K}$ of the group intertwined by an element from a maximal Abelian subgroup $\mathcal{A}$: $G=KAK^{\prime},\,K,K^{\prime}\in\mathcal{K},\,A\in\mathcal{A}$. Such idea of non-compact averaging has been introduced in [36, 8] in the context of averaging quantum states over SLOCC-type operations, and generalized to averaging over arbitrary finite-dimensional matrix representations of reductive groups in the recent work [28]. Following [28], let us take arbitrary matrix representation $\pi_{\mathcal{G}}$ of the group $\mathcal{G}$ and define the averaging of a quantum state with respect to the chosen representation as follows:
$\mathcal{M}_{\mathcal{G}}(\rho)=\int_{\mathcal{K}\times\mathcal{A}\times\mathcal{K}}\left(\pi_{\mathcal{G}}(K)\pi_{\mathcal{G}}(A_{\textrm{n}})\pi_{\mathcal{G}}(K^{\prime})\right)\rho\left(\pi_{\mathcal{G}}(K)\pi_{\mathcal{G}}(A_{\textrm{n}})\pi_{\mathcal{G}}(K^{\prime})\right)^{\dagger}\operatorname{d}\!K\operatorname{d}\!A\operatorname{d}\!K^{\prime}.$ (22)
in which the integration is defined over the Cartan “KAK” decomposition of $\mathcal{G}$. The integration over the two copies of maximally compact components $\mathcal{K},\mathcal{K}^{\prime}$ is performed uniformly, which means that $\int_{\mathcal{K}}\operatorname{d}\!K$ represents Haar integral over $\mathcal{K}$. Integration over $\mathcal{A}$ is performed with respect to any finite normalized measure on this (in general non-compact) manifold, but there appears one crucial assumption: elements of $\mathcal{A}$ enter the integral as normalized matrices: $AA_{\textrm{n}}$=A/||A||, in order to guarantee trace nonincreasing character of the map (22).
Our aim is to define twirling of quantum channels with respect to the representation $\pi_{\mathcal{G}}$ applied as pre- and post-processing operation:
$\mathcal{T}_{G}^{\operatorname{in}}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)^{\dagger},\,\,\sigma_{\operatorname{in}}\in\mathcal{B}(\mathcal{H}_{\operatorname{in}})$
$\mathcal{T}_{G}^{\operatorname{out}}(\sigma_{\operatorname{out}})$ $=$ $\pi^{\operatorname{out}}_{{\mathcal{G}}}(G)\sigma_{\operatorname{out}}\pi^{\operatorname{out}}_{{\mathcal{G}}}(G)^{\dagger},\,\,\sigma_{\operatorname{out}}\in\mathcal{B}(\mathcal{H}_{\operatorname{out}}).$ (23)
In order to follow the analogy with (4) we have to define the counterpart of the $\mathcal{T}_{U^{-1}}$ operation from definition (4). Formally every element $G$ from the group $\mathcal{G}$ is invertible, however, the demand of trace-nonincreasing character of a quantum channel forces us to integrate over normalized elements from the group $\mathcal{G}$ in the twirling channel (22). Now, in general normalized matrices, which represent trace-nonincreasing quantum channels, are not invertible as quantum operations. Therefore we define twirling of quantum channels with respect to $\mathcal{G}$ operations as follows:
$\mathcal{M}^{\Phi}_{\mathcal{G}}(\rho)=\int\operatorname{d}\!G\,\mathcal{T}_{G}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{G}^{\operatorname{in}\dagger}(\rho),$ (24)
in which we define:
$\mathcal{T}_{G}^{\operatorname{in}\dagger}(\sigma_{\operatorname{in}})$ $=$ $\pi^{\operatorname{in}}_{{\mathcal{G}}}(G)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{{\mathcal{G}}}(G).$ (25)
Note that the integration in (24) is understood in the sense of Cartan integral (22). In the case of a compact group $\mathcal{G}$ with trivial $\mathcal{A}$ this definition is fully equivalent to (4).
In analogy with (2) the Choi matrix of a channel twirled with respect to a group $\mathcal{G}$ reads:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\left(\mathcal{M}^{\Phi}_{\mathcal{G}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}$
$=$ $\left(\int\operatorname{d}\!G(\mathcal{T}_{G}^{\operatorname{out}}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\circ(\mathcal{T}_{G}^{\operatorname{in}\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\right)\rho_{\operatorname{ME}}$
$=$ $\int\operatorname{d}\!G\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\left((\Phi\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}})\left[\left(\pi_{\mathcal{G}}^{\operatorname{in}}(G)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\rho_{\operatorname{ME}}\left(\pi_{\mathcal{G}}^{\operatorname{in}}(G)\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right)\right]\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)^{\dagger}\otimes{\rm 1\kern-2.20001ptl}_{\operatorname{in}}\right).$
Now all of the derivations (6)–(2) concerning the unitary case hold, since they do not utilize the property of unitarity of the underlying representations, and we have the following expression for the Choi matrix of the channel (24):
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int\operatorname{d}\!G\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)\right)J_{\Phi}\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)\right)^{\dagger}.$ (27)
Due to distributivity of the tensor product with respect to ordinary matrix product, we have the following Cartan decompositon:
$\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(G)=\left(\pi_{\mathcal{G}}^{\operatorname{out}}(K)\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(K)\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(A_{\textrm{n}})\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(A_{\textrm{n}})\right)\left(\pi_{\mathcal{G}}^{\operatorname{out}}(K^{\prime})\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}(K^{\prime})\right),$ (28)
in which we omit bar in the middle term since $A_{\textrm{n}}$ is a real diagonal matrix. Since Cartan decomposition commutes with taking representations of the components [28], it can be seen that $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}\equiv\pi_{\mathcal{K}}$, when acting solely on compact elements $K\in\mathcal{K}$ of the group $\mathcal{G}$, is itself a representation of the maximal compact subgroup $\mathcal{K}$ of $\mathcal{G}$. Analogously, in the same sense $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{A}}\equiv\pi_{\mathcal{A}}$, when acting solely on non-compact elements $A\in\mathcal{A}$ can be treated as a representation of the maximal Abelian subgroup $\mathcal{A}$ of $\mathcal{G}$. Hence the integral (27) becomes:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int_{\mathcal{K}\times\mathcal{A}\times\mathcal{K}}\left(\pi_{\mathcal{K}}(K)\pi_{\mathcal{A}}(A_{\textrm{n}})\pi_{\mathcal{K}}(K^{\prime})\right)J_{\Phi}\left(\pi_{\mathcal{K}}(K)\pi_{\mathcal{A}}(A_{\textrm{n}})\pi_{\mathcal{K}}(K^{\prime})\right)^{\dagger}\operatorname{d}\!K\operatorname{d}\!A\operatorname{d}\!K^{\prime}.$ (29)
As shown in [8] such integral decomposes into a product of compact twirling of the input operator and of the square of the non-compact part:
$\mathcal{M}_{\mathcal{G}}(\rho)=\mathcal{M}_{\mathcal{K}}(\rho)\mathcal{M}_{\mathcal{K}}\left(\pi_{\mathcal{A}}(A_{\textrm{n}})^{2}\right),$ (30)
which can be directly represented as:
$\mathcal{M}_{\mathcal{G}}(\rho)=\sum_{k}\left(\frac{\beta_{k}}{D^{k}}\right)\mathcal{M}_{\mathcal{K}}^{(k)}(\rho)=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(\rho\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}],$ (31)
where the coefficients $\beta_{k}$ are integrals over the non-compact components:
[…]
and the operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}]$ are Schur operator basis found by applying generalized duality theorem to the representation $\pi_{\mathcal{K}}=\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$. Therefore, we finally have:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}]\right)\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}].$ (33)
In full analogy with the previous case one may want to avoid using contragradient representations in the generalized Schur-Weyl duality [^3] Note that although formally $\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}$ is not a contragradient representation to $\pi_{\mathcal{G}}^{\operatorname{in}}$, for a non-unitary representation $\pi_{\mathcal{G}}^{\operatorname{in}}$, the restriction $\bar{\pi}_{\mathcal{K}}^{\operatorname{in}}$ is contragradient to $\pi_{\mathcal{K}}^{\operatorname{in}}$, since we can without loss of generality always take unitary $\pi_{\mathcal{K}}^{\operatorname{in}}$.. Utilizing Lemma 3 the formula (27) becomes:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int\operatorname{d}\!G\left[\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(G)\right)J_{\Phi}^{\Gamma}\left(\pi_{\mathcal{G}}^{\operatorname{out}}(G)\otimes\pi_{\mathcal{G}}^{\operatorname{in}}(G)\right)^{\dagger}\right]^{\Gamma},$ (34)
which can be analogously, using Cartan decomposition, decomposed into:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]\right)^{\Gamma},$ (35)
where now the operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}]$ are elements of the Schur operator basis related with decomposing the representations of $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}=\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\pi_{\mathcal{K}}^{\operatorname{in}}$, which are typically much easier to construct. Indeed, let us take as an example the case of $\mathcal{G}=\textrm{SL}(d,\mathbb{C})$, for which the maximally compact Cartan component $\mathcal{K}$ is the special unitary group $\textrm{SU}(d)$. In full analogy to the unitary case we assume collective action of group elements on input and output spaces:
$\pi_{\textrm{SL}}^{\operatorname{in}}(L)$ $=$ $L^{\otimes t_{\operatorname{in}}},\,\,L\in\textrm{SL}(d,\mathbb{C}),$
$\pi_{\textrm{SL}}^{\operatorname{out}}(L)$ $=$ $L^{\otimes t_{\operatorname{out}}},\,\,L\in\textrm{SL}(d,\mathbb{C}).$ (36)
In this case the Choi matrix of a twirled channel reads after (34):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}$ $=$ $\int_{\textrm{SL}(d,\mathbb{C})}\operatorname{d}\!L\left[\left(L^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(L^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}.$ (37)
Now the restricted representation $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\pi_{\mathcal{G}}^{\operatorname{in}}|_{\mathcal{K}}$ reads in this case simply $U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, for $U\in\textrm{SU}(d)$, and we have, following (35):
$J_{\mathcal{M}^{\Phi}_{\textrm{SL}}}=\sum_{k}\frac{1}{D^{k}_{U}}\left(\frac{\beta_{k}}{D^{k}}\right)\sum_{\lambda_{1}\lambda_{2}=1}^{D^{k}_{C}}\operatorname{Tr}\left(J_{\Phi}^{\Gamma}\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}\dagger}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)\left(\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]\right)^{\Gamma},$ (38)
where now $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}[U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}]$ are ordinary permutation operators corresponding to standard Schur-Weyl duality for $\textrm{U}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$ and the permutation group $\textrm{S}_{(t_{\operatorname{in}}+t_{\operatorname{out}})}$.
4 Dual averaging of channels
In this section we utilize the concept of generalized finite unitary averaging introduced in a recent work [28] in the context of averaging of quantum channels. The idea of generalized finite unitary averaging, originally introduced for averaging quantum states over general symmetry operations, also bases on generalized duality for group representations, however, in analogy to the concept of unitary $t$-designs, it keeps the averaging procedure on the original group representation side, and not on its commutant, as in formulas (14) and (33), in which the Choi matrix of the twirled channel is expressed in terms of operators $\hat{\Lambda}_{k}^{\lambda_{1}\lambda_{2}}$, which span the commutant of tensor product of the original symmetry group. Indeed, the idea in [28] is to represent the Cartan-decomposition-based average of a quantum state with respect to arbitrary symmetry group (22) as a probabilistic mixture of unitary-$1$-designs on invariant subspaces:
$\mathcal{M}_{\mathcal{G}}(\rho)=\sum_{k\in\mathcal{I}}p_{k}\mathcal{D}_{\mathcal{K}}^{(k)}(\rho),$ (39)
in which $\mathcal{D}_{\mathcal{K}}^{(k)}$ represents unitary-$1$-design on $D^{k}_{G}$-dimensional subspace irreducible under the action of the restricted representation $\pi_{\mathcal{K}}$, for $\mathcal{K}$ being the maximal compact subgroup of $\mathcal{G}$:
$\mathcal{D}_{\mathcal{K}}^{(k)}(\rho)=\frac{1}{\left(D^{k}_{G}\right)^{2}}\sum_{l=1}^{\left(D^{k}_{G}\right)^{2}}\tilde{\gamma}_{l}^{(k)}\rho\,\tilde{\gamma}_{l}^{(k)\dagger},$ (40)
and the probabilities $p_{k}$ are specified in full analogy to (31) as $p_{k}=\frac{\beta_{k}}{D^{k}}$. Unitary-$1$-designs are specified by choosing a unitary operator basis $\{\gamma^{(k)}_{l}\}_{l=1}^{\left(D^{k}_{G}\right)^{2}}$ on space $\mathbb{C}^{D^{k}_{G}}$, and further they are extended to act on the entire representation space using basis operators $\hat{\Pi}_{k}^{m_{1}m_{2}}[\pi_{\mathcal{K}}]$ (2) specified by generalized group duality:
$\tilde{\gamma}^{(k)}_{l}=\sum_{m_{1},m_{2}=1}^{D^{k}_{G}}{\left(\gamma^{(k)}_{l}\right)}_{m_{1}m_{2}}\hat{\Pi}^{m_{1}m_{2}}_{k}[\pi_{\mathcal{K}}].$ (41)
Utilizing the above formalism one can express the Choi matrix $J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ of a twirled channel (27) in a design-like form:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k\in\mathcal{I}}\left(\frac{\beta_{k}}{D^{k}}\right)\mathcal{D}_{\mathcal{K}}^{(k)}(J_{\Phi})=\sum_{k\in\mathcal{I}}\left(\frac{\beta_{k}}{D^{k}}\right)\frac{1}{\left(D^{k}_{G}\right)^{2}}\sum_{l=1}^{\left(D^{k}_{G}\right)^{2}}\tilde{\gamma}_{l}^{(k)}J_{\Phi}\,\tilde{\gamma}_{l}^{(k)\dagger},$ (42)
where the operators (41) are defined for $\hat{\Pi}^{m_{1}m_{2}}_{k}[\pi_{\mathcal{K}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{K}}^{\operatorname{in}}]$. In principle in the above formulas one can use arbitrary unitary operator basis, however the typical choice due to its simplicity of construction is the Heisenberg-Weyl unitary operator bases [37], comprising operators:
$\{\omega_{D^{k}_{G}}^{ij}Z_{D^{k}_{G}}^{i}X_{D^{k}_{G}}^{j}\}_{i,j=0}^{{D^{k}_{G}}-1}$
in which the complex root of unity reads: $\omega_{{D^{k}_{G}}}=e^{2i\pi/{D^{k}_{G}}}$, and the operators $Z,X$ are unitary generalizations of Pauli matrices $\sigma_{z},\sigma_{x}$.
5 Channel $t$-designs
The methods of averaging quantum channels described in previous sections are based on generalized group duality and therefore, though finite, they demand performing projections on irreducible subspaces related with representations of either the original symmetry group for input and output ports or the commutant of its action. One may expect a finite averaging method which respects the operator structure of the original group actions, as in the case of unitary $t$-designs. We will provide such a generalization utilizing Lemma 3 for the following case:
$\pi_{\mathcal{G}}^{\operatorname{in}}(G)$ $=$ $G^{\otimes t_{\operatorname{in}}},\,\,G\in\mathcal{G},$
$\pi_{\mathcal{G}}^{\operatorname{out}}(G)$ $=$ $G^{\otimes t_{\operatorname{out}}},\,\,G\in\mathcal{G},$ (43)
in which $\mathcal{G}$ is an arbitrary subgroup of $\textrm{GL}(d,\mathbb{C})$. We utilize the following lemma:
Lemma 4.
Let $\{G_{i},w_{i}\}$ be a (weighted) group $t$-design, namely:
$\int_{\mathcal{G}}\operatorname{d}\!G\,G^{\otimes t}\rho\,G^{\otimes t\dagger}=\sum_{i\in\mathcal{X}}w_{i}G_{i}^{\otimes t}\rho\,G_{i}^{\otimes t\dagger}.$ (44)
Then the following holds:
$\int_{\mathcal{G}}\operatorname{d}\!G\,\left[G^{\otimes t}\rho^{\Gamma}\,G^{\otimes t\dagger}\right]^{\Gamma}=\sum_{i\in\mathcal{X}}w_{i}\left[G_{i}^{\otimes t}\rho^{\Gamma}\,G_{i}^{\otimes t\dagger}\right]^{\Gamma}.$ (45)
Proof.
The above holds due to: (i) linearity of partial transpose, (ii) the fact that (44) holds for any $\rho$ (in fact $\rho$ is redundant in this formula). ∎
Utilizing the above lemma we can define channel $t$-design, where $t=t_{\operatorname{in}}+t_{\operatorname{out}}$ as a finite set of matrices and weights $\{G_{i},w_{i}\}$ which reconstruct the Choi matrix of an averaged channel:
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}$ $=$ $\int_{\mathcal{G}}\operatorname{d}\!G\left[\left(G^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(G^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}$ (46)
$=$ $\sum_{i\in\mathcal{X}}w_{i}\left[\left(G_{i}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)J_{\Phi}^{\Gamma}\left(G_{i}^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\right)^{\dagger}\right]^{\Gamma}.$
The above construction is very general and can be applied whenever one can construct $t$-design for a given symmetry group $\mathcal{G}$ and demanded value of $t=t_{\operatorname{in}}+t_{\operatorname{out}}$. In the case of $\mathcal{G}$ being a unitary group there is a huge literature on this topic [38, 39, 40, 41, 42, 7], whereas in the case of $\mathcal{G}=\textrm{SL}(2,\mathbb{C})$ one can construct respective designs as product designs comprising unitary ones and Gauss-type quadratues, see [36] for detailed presentation.
Relation to designs for channels.
Several notions of designs for channels have been proposed in the literature, often by pushing forward a design on a larger unitary group through a fixed Stinespring dilation or by imposing moment-matching conditions on ensembles of CPTP maps, see e.g. [43]. The channel $t$-design notion in Eq. (43) is tailored to the present twirling problem: it is the minimal finite condition required to reproduce the averaged Choi operator for collective pre/post actions, and it follows directly from the partial-transpose trick and any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$. This perspective complements other channel-design frameworks by making the symmetry and the induced representation on the Choi space explicit.
6 Conclusions and outlook
We have shown that channel twirling is most naturally and generally formulated at the level of Choi operators, where it becomes a genuine group twirl under the induced output–input representation. From this perspective, channel averaging is no longer merely an operational pre- and post-processing prescription, but a representation-theoretic projection acting directly on the Choi operator space. This shift of viewpoint provides a unified and explicit description of channel twirling that applies to arbitrary input and output representations and does not rely on case-by-case constructions.
While symmetry-restricted quantum channels—covariant or equivariant—have been extensively studied, existing works typically focus on characterizing the set of invariant channels rather than on the twirling map itself. In contrast, the present work treats the twirl as an explicit linear superoperator acting on Choi operators. Concretely, the Haar (or Cartan) average defines a completely positive linear map whose range coincides with the commutant of the induced representation, so that the twirl realizes a projection onto this invariant subspace in the operator-algebraic sense. This distinction is crucial: instead of merely identifying the structural constraints that invariant channels must satisfy, we obtain a constructive formula for the averaging map itself, allowing one to compute the averaged channel directly rather than inferring its form indirectly from symmetry considerations.
In the collective setting, where the symmetry acts as tensor powers on multiple input and output subsystems, the relevant commutant is the walled Brauer algebra. A central insight of this work is that the technical complexity associated with mixed Schur–Weyl duality can be bypassed by a partial-transpose reduction: the channel twirl is mapped to an ordinary Schur–Weyl twirl of the partially transposed Choi operator under a single tensor-power representation. This reduction replaces contragredient actions by a final partial transpose and enables practical evaluation using standard permutation-operator techniques, without constructing explicit walled Brauer idempotents or mixed Schur transforms.
We further extended the Choi-level formulation of channel twirling beyond compact symmetries acting via unitary representations to reductive, generally non-compact groups via Cartan decomposition. In this setting, where uniform averaging is not available, the twirl decomposes into invariant-sector projections determined by the maximal compact subgroup, weighted by coefficients that depend solely on the Abelian Cartan component. To our knowledge, this provides the first systematic and constructive treatment of channel twirling for non-compact transformations at the Choi level, clarifying both its algebraic structure and its normalization properties.
Two finite realizations of channel averaging were also developed. First, we showed that the Cartan twirl admits a dual representation as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors through embedded operator bases. Second, motivated by unitary designs, we introduced a notion of channel $t$-design tailored to the present setting: any weighted group $t$-design with $t=t_{\mathrm{in}}+t_{\mathrm{out}}$ induces an exact finite reconstruction of the averaged Choi operator. Together, these constructions demonstrate that the abstract Choi-level twirl admits finite, symmetry-respecting implementations without leaving the representation-theoretic framework.
Several directions follow naturally from this work. On the constructive side, it is important to develop efficient exact and approximate finite averaging sets for broader classes of symmetry groups and representations. A key open problem is to quantify the approximation error incurred when replacing Haar or Cartan twirling by a finite weighted sum (design-like implementation). Concretely, one would like bounds on the deviation of the resulting twirling superoperator from the ideal one, either at the Choi level (e.g. in trace or Hilbert–Schmidt norm for Choi operators) or directly at the channel level (e.g. in diamond norm), with explicit dependence on $t_{\mathrm{in}}+t_{\mathrm{out}}$, the local dimension, and the quality/size of the design.
From an algorithmic perspective, the partial-transpose reduction suggests scalable implementations of channel twirls based on permutation operators and Schur–Weyl projectors, with complexity controlled by $t_{\mathrm{in}}+t_{\mathrm{out}}$ rather than by the full mixed Schur–Weyl machinery. More broadly, treating channel symmetries directly at the Choi level appears essential for extending symmetry-based reductions to approximate symmetries, non-identical local dimensions, and learning problems involving structured quantum noise.
Acknowledgements
MM acknowledges support from the National Science Center (NCN), Poland, under Project Opus No. 2024/53/B/ST2/02026. ZP and ŁP acknowledge support from the National Science Center (NCN), Poland, under Project Opus No. 2022/47/B/ST6/02380.